OpenStudy (anonymous):

The width of a rectangle is 6 inches less than twice its length. Find the dimension of the rectangle if its area is 108 square inches.

OpenStudy (triciaal):

width = w, length = l given w = 2l-6 area of a rectangle = l*w =108 L*(2l-6) = 108 solve for l substitute in w = 2l-6 to find w